# 7491, Trondheim, Norway

### v1

@@ 1 Introduction
Studying structure of solutions plays an important role in vector optimization theory, see, for example (lucbook, , Chapter 6) and (ehrgott2005multicriteria, , Chapter 3). The most effective method for such studying is considering solution sets as the images of solution mappings based on scalarization techniques.
It is well-known that the image of a connected set under a continuous single-valued mapping is connected. Naccache in naccache1978connectedness , Warburton in warburton1983quasiconcave , and Hiriart-Urruty in hiriart1985images extended the result for a set-valued mapping having connected values, provided that the mapping is upper (or lower) semi-continuous, see Theorem 2.1. This criterion opened a door to study connectedness of solution sets of not only vector optimization problems but also of set optimization problems, vector variational inequalities, vector equilibrium problems. There are numerous works related to this topic, we can only list a few of them, for example luc1987connectedness ; sun1996connectedness ; gong2007connectedness ; yen2011monotone ; han2018existence ; sharma2022connectedness ; peng2019connectedness ; han2022connectedness ; anh2022connectedness . Almost all of the obtained results require generalized convexity or monotonicity assumptions to ensure connected-valuedness, and require boundedness of the constraint sets or certain coercivity conditions to guarantee semi-continuity of solution mappings.
@@
In the second part, we apply the above results to vector optimization based on two solution mappings associated to two scalarization methods. We obtain new properties related to unboundedness of the (weak) Pareto solution set of a problem in both cases: with a generalized convexity assumption and without imposing generalized convexity. For the second case, in Theorem 5.2 we prove that if the domain of a solution mapping between two aforementioned ones is non-closed then the weak Pareto solution set is unbounded.
For a problem governed by semi-strictly quasi-convex functions, to ensure disconnectedness of all components of the weak Pareto solution set, Hoa, Huy, Phuong and Yen in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite. In Theorem 6.1, we introduce two other criteria for quasi-convex problems. As a result, when the objective functions are strictly quasi-convex, the Pareto solution set and weak Pareto solution set are coincident, and if this set is disconnected then all the components are unbounded.
@@ 2 Preliminaries
Let $X$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We consider set-valued mapping $S$ from $X$ to $\operatorname{\mathbb{R}}^{n}$, $S:X\rightrightarrows\operatorname{\mathbb{R}}^{n}$. The domain of $S$ is the set of all elements $u$ in $X$ such that $S(u)$ is non-empty:
@@
is a closed subset of $X\times\operatorname{\mathbb{R}}^{n}$ then one says that $S$ has closed graph. If $S$ has closed graph then $S(u)$ is a closed set for all $u\in X$. The inverse mapping of $S$, denoted by $S^{-1}:\operatorname{\mathbb{R}}^{n}\rightrightarrows X$, is defined by
@@
The mapping $S$ is locally bounded at $\bar{u}$ if there exists an open neighborhood $U$ of $\bar{u}$ such that $S(U\cap X)$ is bounded. If $S$ is locally bounded at every $u\in X$ then $S$ is said to be locally bounded on $X$. The mapping $S$ is upper semi-continuous at $\bar{u}\in X$ if for any open set $V\subseteq\operatorname{\mathbb{R}}^{n}$ such that $S(\bar{u})\subseteq V$ there exists an open neighborhood $U$ of $\bar{u}$ such that $S(u)\subseteq V$ for all $u\in U$. If $S$ is upper semi-continuous at every $u\in X$ then $S$ is said to be upper semi-continuous on $X$.
@@
Let $Y$ be a nonempty subset of $\operatorname{\mathbb{R}}^{n}$. One says that $Y$ is connected if there do not exist two nonempty disjoint subsets $A,B$ of $Y$ with $Y\subseteq A\cup B$ and two open subsets $U,V$ in $\operatorname{\mathbb{R}}^{n}$ such that $A\subseteq U$, $B\subseteq V$ and $U\cap V=\emptyset$. Otherwise $Y$ is called disconnected. A nonempty subset $A\subseteq Y$ is said to be a connected component of $Y$ if $A$ is connected and it is not a proper subset of any connected subset of $Y$. Recall that when $Y$ is connected, if the set $A\subseteq Y$ is closed and open in $Y$ then $A=Y$. Moreover, if $Y$ is connected then the union $Y\cup\{y\}$, where $y$ belongs to the closure $\operatorname{cl}(Y)$, is also connected.
One says that the set-valued mapping $S$ has connected (convex, bounded) values if $S(x)$ is connected (convex, bounded) for every $x$ on the domain of $S$.
In order to be convenient for readers to compare our developments with the before-mentioned result by Naccache, Warburton, and Hiriart-Urruty, we recall the following theorem; its proofs can be found in (naccache1978connectedness, , Lemma 2.1), (warburton1983quasiconcave, , Theorem 3.1), or (hiriart1985images, , Theorem 3.1):
@@
Assume that $X$ is connected and $S$ has connected values. If $S$ is upper semi-continuous on $X$, then the image $S(X)$ is connected.
@@ 3 Set-valued mappings having non-closed domains
Let $A$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We want to emphasize that, throughout the present section, the set-valued mapping
$\Phi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Phi(a),$
has closed graph. We denote by $\operatorname{ri}A$ its relative interior.
The proofs of the main results in the present and next sections are based on the following lemma:
@@
If $\operatorname{\mathcal{A}}$ is a bounded connected component of $\Phi(A)$, then $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed. Furthermore, if $\Phi(A)$ is a bounded, then $\operatorname{dom}\Phi$ is closed.
@@
Let $\operatorname{\mathcal{A}}$ be a bounded connected component of $\Phi(A)$. To prove that $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed, we suppose that the sequence $\{a^{k}\}\subseteq\Phi^{-1}(\operatorname{\mathcal{A}})$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$. Indeed, let $\{x^{k}\}\subseteq\operatorname{\mathcal{A}}$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$. By the boundedness of $\operatorname{\mathcal{A}}$, we can assume that $x^{k}\to\bar{x}$. Since the graph of $\Phi$ is closed, $(\bar{a},\bar{x})$ belongs to $\operatorname{gph}\Phi$, thus $\bar{x}\in\Phi(\bar{a})$. As $\operatorname{\mathcal{A}}$ is connected, the union $\operatorname{\mathcal{A}}\cup\{\bar{x}\}$ is connected. This means that $\bar{x}\in\operatorname{\mathcal{A}}$. We thus get $\bar{x}\in\operatorname{\mathcal{A}}\cap\Phi(\bar{a})\neq\emptyset$, and then $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$ by definition. Therefore, the inverse image $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed.
The second conclusion is proven similarly. ∎
@@
Assume that $\operatorname{dom}\Phi$ is non-closed. Then, $\Phi(A)$ is unbounded. Moreover, if additionally $A$ is convex and $\Phi(\operatorname{ri}A)$ is nonempty then the latter set is also unbounded.
@@
Now let $A$ be convex. Suppose that $\Phi(\operatorname{ri}A)$ is nonempty but on the contrary the set is bounded. Let $\overline{\Sigma}$ be the closure of the following set:
@@
Clearly, one has $\Sigma\subseteq\operatorname{gph}\Phi$. Since $\operatorname{gph}\Phi$ is closed, the closure of $\Sigma$ is also a subset of $\operatorname{gph}\Phi$, i.e. $\overline{\Sigma}\subseteq\operatorname{gph}\Phi$. Let $\pi:\operatorname{\mathbb{R}}^{m+n}\to\operatorname{\mathbb{R}}^{n}$ be the projection defined by $\pi(t_{1},\dots,t_{m},t_{m+1},\dots,t_{m+n})=(t_{m+1},\dots,t_{m+n})$. Clearly, the image of $\overline{\Sigma}$ under $\pi$ is bounded and
$\Phi(\operatorname{ri}A)\subseteq\pi(\overline{\Sigma})\subseteq\Phi(A).$
We consider a new mapping $\Phi^{\prime}$ defined by
@@
One can check that the graph of $\Phi^{\prime}$ coincides with $\overline{\Sigma}$. Hence, $\Phi^{\prime}$ also has closed graph. By boundedness of $\pi(\overline{\Sigma})$, $\Phi^{\prime}(A)=\pi(\overline{\Sigma})$ is also bounded. Applying the second conclusion in Lemma 2 to $\Phi^{\prime}$, $\operatorname{dom}\Phi^{\prime}$ is closed. Furthermore, it is not difficult to see the following inclusions:
@@
By convexity of $A$ and density of $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ in $\operatorname{dom}\Phi$, $\operatorname{dom}\Phi^{\prime}$ is also dense in $\operatorname{dom}\Phi$. It follows that $\operatorname{dom}\Phi$ is closed. This is a contradiction. ∎
@@
From Example 6 we consider the set-valued mapping that is the solution mapping $\Theta_{1}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{2}$ of the there given vector optimization problem:
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\pm\sqrt{\frac{1-\xi_{1}}{\xi_{1}}},\frac{\xi_{1}}{1-\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ \xi_{1}\in(0,1),\\ \emptyset&\ \ \hbox{ if }\ \xi_{1}\in\{0,1\}.\\ \end{array}\right.$ (3.1)
@@
$\Theta_{1}([0,1])=\big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}^{2}x_{2}=1\big{\}}=\operatorname{\mathcal{A}}\dot{\cup}\operatorname{\mathcal{B}},$ (3.2)
@@ 4 Set-valued mappings having connected values
It is important to underline that, throughout the section, we consider another set-valued mapping
@@
Suppose that $\Psi(A)$ is disconnected and $\operatorname{\mathcal{A}}$ is a connected component of $\Psi(A)$. On the contrary, we assume that $\operatorname{\mathcal{A}}$ is bounded. Let $\operatorname{\mathcal{B}}$ is another connected component of $\Psi(A)$. Then there exists an open set $V$ containing $\operatorname{\mathcal{A}}$ such that $V\cap\operatorname{\mathcal{B}}=\emptyset$. Recall from Lemma 3 that the two sets $\Psi^{-1}(\operatorname{\mathcal{A}})$ and $\Psi^{-1}(\operatorname{\mathcal{B}})$ are disjoint.
By the closedness of the graph of $\Psi$ and Lemma 2, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is closed. We now point out that $\Psi^{-1}(\operatorname{\mathcal{A}})$ is open in $\operatorname{dom}\Psi$. Indeed, let $a$ be an arbitrary point in $\Psi^{-1}(\operatorname{\mathcal{A}})$, then $\emptyset\neq\Psi(a)\subseteq\operatorname{\mathcal{A}}$ is bounded. It follows that $a\in\mathcal{O}$. By the assumption, $\Psi$ is upper semi-continuous at $a$. With the open set $V$ defined above, there is an open set $U$ in $\operatorname{dom}\Psi$ such that $\Psi(a^{\prime})\subseteq V$ for all $a^{\prime}\in U$. Hence, the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ is proved.
From above, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is both open and closed in the connected set $\operatorname{dom}\Psi$. It follows that $\Psi^{-1}(\operatorname{\mathcal{A}})=\operatorname{dom}\Psi$, while $\operatorname{dom}\Psi$ contains $\Psi^{-1}(\operatorname{\mathcal{B}})\neq\emptyset$. This is impossible because of $\Psi^{-1}(\operatorname{\mathcal{A}})\cap\Psi^{-1}(\operatorname{\mathcal{B}})=\emptyset$. Thus, $\operatorname{\mathcal{A}}$ must be unbounded. ∎
To illustrate above theorem, we consider an example coming from linear fractional vector optimization.
@@
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\{(1,0)\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2},\\ [1,+\infty)\times\{0\}&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{1}{2},\\ \left\{\Big{(}\frac{2-3\xi_{1}}{2\xi_{1}-1},2\Big{)}\right\}&\ \ \hbox{ if }\ \frac{1}{2}<\xi_{1}<\frac{2}{3},\\ \{0\}\times[2,+\infty)&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{2}{3},\\ \emptyset&\ \ \hbox{ if }\ \frac{2}{3}<\xi_{1}\leq 1.\\ \end{array}\right.$
The domain of $\Theta_{1}$ is the interval $[0,\frac{2}{3}]$ which is connected. Moreover, the set $\mathcal{O}$ is $[0,\frac{1}{2})\cup(\frac{1}{2},\frac{2}{3})$ and $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Hence, $\Theta_{1}$ satisfies the condition in above theorem. The image has two connected components, both of them are unbounded.
The below theorem proposes another criterion to replace connectedness of $\operatorname{dom}\Psi$ in Theorem 4.1.
@@
The assumption related to openness of $\mathcal{O}$ and local boundedness or upper semi-continuity of $\Psi$ on $\mathcal{O}$ in Theorem 4.2 seems heavily in general. However, it is satisfied for several cases such as solution mappings in strictly quasi-convex vector optimization (see Lemma 6), linear fractional vector optimization, or monotone affine/general vector inequality yen2011monotone ; hieu2021disconnectedness .
@@
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\frac{1}{1-2\xi_{1}},\frac{-1}{1-2\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\ \hbox{ or }\ \frac{1}{2}<\xi_{1}\leq 1,\\ \emptyset&\ \ \hbox{ if }\qquad\qquad\ \,\xi_{1}=\frac{1}{2}.\\ \end{array}\right.$
Clearly, $\operatorname{dom}\Theta_{1}=[0,\frac{1}{2})\cup(\frac{1}{2},1]=\mathcal{O}$ is not connected but open in $[0,1]$. One observers that $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Its image has two connected components, both of them are unbounded.
@@
We consider the mapping given in Example 3. The mapping satisfies the condition in Proposition 1. Clearly, the number of connected components of $\operatorname{dom}\Theta_{1}$ equals the one of the image $\Theta_{1}([0,1])$.

### v2

@@ 1 Introduction
Studying the structure of solution sets plays an important role in vector optimization theory; see, for example (lucbook, , Chapter 6) and (ehrgott2005multicriteria, , Chapter 3). The most effective method for such studies is to consider the solution sets in question as the images of the solution mappings obtained by some scalarization techniques.
It is well-known that the image of a connected set under a continuous single-valued mapping is connected. Naccache in naccache1978connectedness , Warburton in warburton1983quasiconcave , and Hiriart-Urruty in hiriart1985images extended the result for a set-valued mapping having connected values, provided that the mapping is upper (or lower) semi-continuous; see Theorem 2.1. This criterion opened a door for studying connectedness of solution sets of not only vector optimization problems but also set optimization problems, vector variational inequalities and vector equilibrium problems. There are numerous works related to this topic, we can only list a few of them, for example luc1987connectedness ; sun1996connectedness ; gong2007connectedness ; yen2011monotone ; han2018existence ; sharma2023connectedness ; peng2019connectedness ; han2022connectedness ; anh2022connectedness . Almost all therein obtained results require generalized convexity or monotonicity assumptions to ensure connected-valuedness, and require boundedness of the constraint sets or certain coercivity conditions to guarantee semi-continuity of solution mappings.
@@
In the second part, we apply the above results to vector optimization based on two solution mappings associated to two scalarization methods. We obtain new properties related to unboundedness of the (weak) Pareto solution set of a problem in both cases: with a generalized convexity assumption and without imposing generalized convexity. For the second case, in Theorem 5.2 we prove that if the domain of one of the two aforementioned solution mappings (denoted by $\Theta_{1}$ and $\Theta_{2}$; see the definitions in Subsection 5.1) is non-closed then the weak Pareto solution set is unbounded. The reverse implication holds for $\Theta_{1}$ if the problem is convex and this solution map has bounded values, see Theorem 6.3. In Theorem 6.2, we have the same conclusion for $\Theta_{2}$ when the objective functions are quasi-convex and non-negative over the constraint set.
For a problem governed by semi-strictly quasi-convex functions, to ensure disconnectedness of all components of the weak Pareto solution set, Hoa et al. (hoa2007unbounded, , Theorem 3.2) used the condition that the number of components of the set is finite. In Theorem 6.1, we introduce two other criteria for this property of quasi-convex problems. As a result, when the objective functions are strictly quasi-convex, the Pareto solution set coincides with the weak Pareto solution set, and if this set is disconnected then all the components are unbounded.
@@ 2 Preliminaries
Let $X$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We consider a set-valued mapping $S$ from $X$ to $\operatorname{\mathbb{R}}^{n}$, $S:X\rightrightarrows\operatorname{\mathbb{R}}^{n}$. The domain of $S$ is the set of all elements $u$ in $X$ such that $S(u)$ is non-empty:
@@
is a closed subset of $X\times\operatorname{\mathbb{R}}^{n}$, then one says that $S$ has closed graph. If $S$ has closed graph, then $S(u)$ is a closed set for every $u\in X$. The inverse mapping of $S$, denoted by $S^{-1}:\operatorname{\mathbb{R}}^{n}\rightrightarrows X$, is defined by
@@
The mapping $S$ is locally bounded at $\bar{u}$ if there exists an open neighborhood $U$ of $\bar{u}$ such that $S(U\cap X)$ is bounded. If $S$ is locally bounded at every $u\in X$, then $S$ is said to be locally bounded on $X$. The mapping $S$ is upper semi-continuous at $\bar{u}\in X$ if for any open set $V\subseteq\operatorname{\mathbb{R}}^{n}$ such that $S(\bar{u})\subseteq V$ there exists an open neighborhood $U$ of $\bar{u}$ such that $S(u)\subseteq V$ for all $u\in U$. If $S$ is upper semi-continuous at every $u\in X$, then $S$ is said to be upper semi-continuous on $X$.
@@
Let $Y$ be a nonempty subset of $\operatorname{\mathbb{R}}^{n}$. One says that $Y$ is connected if there do not exist two nonempty disjoint subsets $A,B$ of $Y$ with $Y\subseteq A\cup B$ and two open subsets $U,V$ in $\operatorname{\mathbb{R}}^{n}$ such that $A\subseteq U$, $B\subseteq V$ and $(U\cap Y)\cap(V\cap Y)=\emptyset$. Otherwise $Y$ is called disconnected. A nonempty subset $A\subseteq Y$ is said to be a connected component of $Y$ if $A$ is connected and it is not a proper subset of any connected subset of $Y$. Recall that when $Y$ is connected, if the set $A\subseteq Y$ is closed and open in $Y$, then $A=Y$. Moreover, if $Y$ is connected then the union $Y\cup\{y\}$, where $y$ belongs to the closure $\operatorname{cl}(Y)$, is also connected.
One says that the set-valued mapping $S$ has connected (convex, bounded) values if $S(u)$ is connected (convex, bounded) for every $u$ in the domain of $S$.
For the convenience of the readers in comparing our developments with the before-mentioned result by Naccache, Warburton, and Hiriart-Urruty, we recall the following theorem; its proofs can be found in (naccache1978connectedness, , Lemma 2.1), (warburton1983quasiconcave, , Theorem 3.1), or (hiriart1985images, , Theorem 3.1).
@@
Assume that $\operatorname{dom}S$ is connected and $S$ has connected values. If $S$ is upper semi-continuous on $X$, then the image $S(X)$ is connected.
@@ 3 Set-valued mappings having non-closed domains
Let $A$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We denote by $\operatorname{ri}A$ its relative interior. Throughout the present section, it is assumed that the following set-valued mapping has closed graph:
$\Phi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Phi(a).$
The proofs of the main results in the present and next sections are based on the following lemma.
@@
If $\operatorname{\mathcal{A}}$ is a bounded connected component of $\Phi(A)$, then $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed. Furthermore, if $\Phi(A)$ is bounded, then $\operatorname{dom}\Phi$ is closed.
@@
Let $\operatorname{\mathcal{A}}$ be a bounded connected component of $\Phi(A)$. To prove that $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed, we suppose that a sequence $\{a^{k}\}\subseteq\Phi^{-1}(\operatorname{\mathcal{A}})$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$. Let $\{x^{k}\}\subseteq\operatorname{\mathcal{A}}$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$. By the boundedness of $\operatorname{\mathcal{A}}$, we can assume that $x^{k}\to\bar{x}$. Since the graph of $\Phi$ is closed, $(\bar{a},\bar{x})$ belongs to $\operatorname{gph}\Phi$, thus $\bar{x}\in\Phi(\bar{a})$. As $\operatorname{\mathcal{A}}$ is connected, the union $\operatorname{\mathcal{A}}\cup\{\bar{x}\}$ is connected. This means that $\bar{x}\in\operatorname{\mathcal{A}}$. We thus get $\bar{x}\in\operatorname{\mathcal{A}}\cap\Phi(\bar{a})$, and then $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$ by definition. Therefore, the inverse image $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed.
We now prove the second conclusion. Suppose that $\Phi(A)$ is bounded. To prove closedness of $\operatorname{dom}\Phi$, we suppose that a sequence $\{a^{k}\}\subseteq\operatorname{dom}\Phi$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\operatorname{dom}\Phi$. Let $\{x^{k}\}\subseteq\Phi(A)$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$. By the boundedness of $\Phi(A)$, we can assume that $x^{k}\to\bar{x}$. Since the graph of $\Phi$ is closed, $(\bar{a},\bar{x})$ belongs to $\operatorname{gph}\Phi$, thus $\bar{x}\in\Phi(\bar{a})$. This implies that $\bar{a}$ belongs to $\operatorname{dom}\Phi$. ∎
@@
Assume that $\operatorname{dom}\Phi$ is non-closed. Then, $\Phi(A)$ is unbounded. Moreover, if additionally $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ is dense in $\operatorname{dom}\Phi$, then $\Phi(\operatorname{ri}A)$ is unbounded.
@@
Let $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ be dense in $\operatorname{dom}\Phi$. Suppose on the contrary that $\Phi(\operatorname{ri}A)$ is bounded. Let $\overline{\Sigma}$ be the closure of the following set:
@@
Clearly, one has $\Sigma\subseteq\operatorname{gph}\Phi$. Since $\operatorname{gph}\Phi$ is closed, the closure of $\Sigma$ is also a subset of $\operatorname{gph}\Phi$, i.e. $\overline{\Sigma}\subseteq\operatorname{gph}\Phi$. Let $\pi:\operatorname{\mathbb{R}}^{m+n}\to\operatorname{\mathbb{R}}^{n}$ be the projection defined by $\pi(t_{1},\dots,t_{m},t_{m+1},\dots,t_{m+n})=(t_{m+1},\dots,t_{m+n})$. It is not difficult to see that
$\Phi(\operatorname{ri}A)\subseteq\pi(\overline{\Sigma})\subseteq\overline{\pi(\Sigma)}.$
Moreover, by the boundedness of $\Phi(\operatorname{ri}A)$ and the definition of $\Sigma$, we see that $\pi(\Sigma)$ is bounded, hence $\overline{\pi(\Sigma)}$ is as well. Thus, we claim that the image of $\overline{\Sigma}$ under $\pi$ is bounded.
We now consider a new mapping $\Phi^{\prime}$ defined by
@@
One can check that the graph of $\Phi^{\prime}$ coincides with $\overline{\Sigma}$. Hence, $\Phi^{\prime}$ also has closed graph. By boundedness of $\pi(\overline{\Sigma})$, $\Phi^{\prime}(A)=\pi(\overline{\Sigma})$ is also bounded. Applying the second conclusion in Lemma 2 to $\Phi^{\prime}$, we get that $\operatorname{dom}\Phi^{\prime}$ is closed. Furthermore, it is not difficult to see the following inclusions:
@@
By the density of $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ in $\operatorname{dom}\Phi$, the above inclusions imply that $\operatorname{dom}\Phi^{\prime}$ is also dense in $\operatorname{dom}\Phi$. It follows that $\operatorname{dom}\Phi$ is closed. This is a contradiction. ∎
@@
From Example 6 we consider the set-valued mapping that is the solution mapping $\Theta_{1}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{2}$ of the vector optimization problem given there:
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\pm\sqrt{\frac{1-\xi_{1}}{\xi_{1}}},\frac{\xi_{1}}{1-\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ \xi_{1}\in(0,1),\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \emptyset&\ \ \hbox{ if }\ \xi_{1}\in\{0,1\}.\\ \end{array}\right.$ (3.1)
@@
$\Theta_{1}([0,1])=\big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}^{2}x_{2}=1\big{\}}=\operatorname{\mathcal{A}}\cup\operatorname{\mathcal{B}},$ (3.2)
@@ 4 Set-valued mappings having connected values
Throughout the section, we consider another set-valued mapping
@@
Suppose that $\Psi(A)$ is disconnected and $\operatorname{\mathcal{A}}$ is a connected component of $\Psi(A)$. On the contrary, we assume that $\operatorname{\mathcal{A}}$ is bounded. Let $\operatorname{\mathcal{B}}$ be another connected component of $\Psi(A)$. Then, there exists an open set $V$ containing $\operatorname{\mathcal{A}}$ such that $V\cap\operatorname{\mathcal{B}}=\emptyset$. Recall from Lemma 3 that the two sets $\Psi^{-1}(\operatorname{\mathcal{A}})$ and $\Psi^{-1}(\operatorname{\mathcal{B}})$ are disjoint.
By the closedness of the graph of $\Psi$ and Lemma 2, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is closed. We now point out that $\Psi^{-1}(\operatorname{\mathcal{A}})$ is open in $\operatorname{dom}\Psi$. Indeed, let $a$ be an arbitrary point in $\Psi^{-1}(\operatorname{\mathcal{A}})$, then $\emptyset\neq\Psi(a)\subseteq\operatorname{\mathcal{A}}$ is bounded. It follows that $a\in\mathcal{O}$. By assumption, $\Psi$ is upper semi-continuous at $a$. With the open set $V$ defined above, there is an open set $U$ in $\operatorname{dom}\Psi$ such that $\Psi(a^{\prime})\subseteq V$ for all $a^{\prime}\in U$. Hence, the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ in $\operatorname{dom}\Psi$ is proved.
From the above, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is both open and closed in the connected set $\operatorname{dom}\Psi$. It follows that $\Psi^{-1}(\operatorname{\mathcal{A}})=\operatorname{dom}\Psi$, while $\operatorname{dom}\Psi$ contains $\Psi^{-1}(\operatorname{\mathcal{B}})\neq\emptyset$. This is impossible because of $\Psi^{-1}(\operatorname{\mathcal{A}})\cap\Psi^{-1}(\operatorname{\mathcal{B}})=\emptyset$. Thus, $\operatorname{\mathcal{A}}$ must be unbounded. ∎
To illustrate Theorem 4.1, we consider an example coming from linear fractional vector optimization.
@@
[…]
The domain of $\Theta_{1}$ is the interval $[0,\frac{2}{3}]$ which is connected. Moreover, the set $\mathcal{O}$ is $[0,\frac{1}{2})\cup(\frac{1}{2},\frac{2}{3})$ and $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Hence, $\Theta_{1}$ satisfies the condition in Theorem 4.1. The image has two connected components, both of them are unbounded.
The below theorem proposes another criterion to replace the connectedness of $\operatorname{dom}\Psi$ in Theorem 4.1.
@@
The assumption related to the openness of $\mathcal{O}$ and local boundedness or upper semi-continuity of $\Psi$ on $\mathcal{O}$ in Theorem 4.2 seems heavy in general. However, it is satisfied for several cases such as solution mappings in strictly quasi-convex vector optimization (see Lemma 6), linear fractional vector optimization, or monotone affine/general vector inequalities yen2011monotone ; hieu2021disconnectedness .
@@
[…]
Clearly, $\operatorname{dom}\Theta_{1}=[0,\frac{1}{2})\cup(\frac{1}{2},1]=\mathcal{O}$ is not connected but open in $[0,1]$. One observers that $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Its image has two connected components, both are unbounded.
@@
We consider the mapping given in Example 3, which satisfies the condition in Proposition 1. Clearly, the number of connected components of $\operatorname{dom}\Theta_{1}$ equals the one of the image $\Theta_{1}([0,1])$.